A train moving at a speed of 40m/s sounds its whistle, which has a frequency of 500Hz. Determine the frequency heard by a stationary observer as the train approaches the observer. The ambient temperature is 24°C.
Answer #1

- U= 40m/s, to = 500 Hz 40=0, = 343 mls Since f= to (Vt 3 f= 500 ( 34 BD It = 566 Hz /